OpenStudy (raffle_snaffle):

Find an equation of the largest sphere with center (5, 4, 9) that is contained in the first octant? Describe process.

OpenStudy (psymon):

I'm new to this stuff too, so someone will correct me if I'm wrong xD If it is in the first octant, then the maximum radius we can have without going outside of the 1st octant would be 4, since our y-coordinate is only 4 units away. So that being said, we just need to know that the equation for a sphere is \[(x-x _{0})^{2}+(y-y _{0})^{2}+(z-z _{0})^{2}=r ^{2}\] So just filling in the blanks, we could just plug in each coordinate into the appropriate part of the equation and also know that the radius of 4 squared will be 16. So this would then give us \[(x-5)^{2}+(y-4)^{2}+(z-9)^{2} = 16\]

OpenStudy (anonymous):

r = 4, center at given point...

OpenStudy (raffle_snaffle):

how do you know the radius r=4?

OpenStudy (psymon):

Because the conditions say that it has to be in the first octant. Any radius greater than that will push the edges of the sphere into another octant.

OpenStudy (raffle_snaffle):

I am thinking right now.

OpenStudy (anonymous):

suppose r > 4, and x = z = 0

OpenStudy (raffle_snaffle):

The first octant?

OpenStudy (psymon):

Well, the question said contained in the first octant O.o

OpenStudy (anonymous):

3 dimensions... each has a + & - direction => 2^3 = 8 octants

OpenStudy (anonymous):

like 2d has 4 quadrants

OpenStudy (psymon):